OpenStudy (anonymous):

Let f(x)= x^2 - 2x + 3. Find F(x+h) - F(x) / h

OpenStudy (anonymous):

\[\frac{(x+h)^2-2(x+h)+3-x^2+2x-3}{h}\] \[\frac{x^2+2xh+h^2-2x-2h+3-x^2+2x-3}{h}\] \[\frac{2xh+h^2-2h}{h}=\frac{h(2x+h-2)}{h}=2x+h-2\]